OpenStudy (tt0860):

-5 + 10x - 7y + 10 - 3y + 5x a. -5x + 5y - 5 2x + 9y - 7x + 5 - 4y - 10 b. = -5x + 15y + 5 5x + 10y - 10x + 5y - 15 + 20 c. = 15x - 10y + 5 5x + 5y - 5 d. 5x + 5y - 5

OpenStudy (mhchen):

Do you know how to combine like-terms? 1 + 1 = 2 a + a = 2a a + b = a + b